OpenStudy (anonymous):

Assume that \sin(x) equals its Maclaurin series for all x. Use the Maclaurin series for sin(6 x^2) to evaluate the integral (integral [0, 0.78]) sin(6 x^2)dx . Your answer will be an infinite series. Use the first two terms to estimate its value.

OpenStudy (anonymous):

\[\int\limits_{0}^{0.78}\sin(6x^2)dx\]

OpenStudy (anonymous):

@mathstudent55

OpenStudy (irishboy123):

writing the Maclaurin sin expansion in terms of t: \(\large sin(t) = \Sigma_{0}^{\infty} \frac{(-1)^n }{(2n+1)! }\ t^{2n+1}\) it seems they are happy for you to stuff \(t = 6 x^2\) in and go from there to use an actual Mac series for \(sin(6 x^2)\), you'd need to go \(f(0) = 0\) \(f'(x) = (sin(6 x^2))' = ...\) and \(f'(0) = ....\) so \(x \ f'(0) = ...\)

OpenStudy (anonymous):

Just plug it in directly?

OpenStudy (anonymous):

What do you think @sourwing

OpenStudy (irishboy123):

"Just plug it in directly?" yes, that is what the question mandates for #1, plug it in term by terms and come up with a nice summation with a \(\Sigma\) in it! for #2, generate a few terms and hope you get the first two quickly. [if you have a crack at the proper Mac expansion for that function, you will be busy for a while :p]

OpenStudy (anonymous):

So basically \[\sum_{n = 0}^{\infty} \frac{ -1^n }{ (2n+1)! } * (6x^2)^{2n+1}\], and solve for n = 0 and n=1?

OpenStudy (irishboy123):

sounds good

OpenStudy (anonymous):

Well what do I do to estimate the value of the series?

OpenStudy (anonymous):

I ask because I still have an x term in my result

OpenStudy (irishboy123):

if you have an 'x' term in your result, well that is a good thing. :p

OpenStudy (anonymous):

Okay lol. Sorry I'm clueless on the procedure here. Should I be evaluating the limit at both of the solutions? They are N_0 = -6x^2 and N_1 = -216x^2

OpenStudy (irishboy123):

now, forget about the \(\Sigma\) and just evaluate the first 2 terms you are not summing, you are just using the first 2 terms of the series make sense?!?! you need x terms because you have to integrate

OpenStudy (irishboy123):

nothing to do with limits, either. just knock out the first 2 terms of the [bogus] expansion

OpenStudy (anonymous):

So get the first two terms in the series, but how do I use that to estimate the value of the series?

OpenStudy (irishboy123):